Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $120 per unit.Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 If the variablRequired 1 Required 2 Required 3 Required 4 At present, the company is selling 14,000 stoves per month. The sales manager isComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Refer to the d
Answer #1

1.

Break-even point in unit sales 4200
Break-even point in dollar sales 504000

Break-even point in unit sales = Total fixed expenses / Contribution per unit = $151200 / ($120 - $84) = $151200/$36 = 4200 units

Break-even point in dollar sales = Total fixed expenses / Contribution margin ratio = $151200/30% = $504000

Contribution margin ratio = Contribution / Sales = ($120 - $84)/$120 = $36/$120 = 30%

OR

Break-even point in dollar sales = 4200 units x $120 = $504000

2. Higher break-even point

If the variable expenses as a percentage of selling price increase, it will lead to a reduction in the contribution margin and hence will result in a higher break-even point.

3.

Outback Outfitters
Contribution Income Statement Present Proposed
14000 Stoves 17500 Stoves
Total Per unit Total Per unit
Sales 1680000 120 1890000 108
Variable costs 1176000 84 1470000 84
Contribution margin 504000 36 420000 24
Fixed costs 151200 151200
Net operating income $ 352800 268800

Proposed units sold = 14000 x 1.25 = 17500 stoves

Proposed selling price per unit = $120 - (10% x $120) = $120 - $12 = $108

Note: The variable cost per unit and monthly fixed costs have been considered to remain the same under the proposed change.

4.

Unit sales needed to attain the target profit 9425

Unit sales for desired net operating income = (Fixed costs + Desired net operating income) / Contribution per unit = ($151200 + $75000)/$24 = $226200/$24 = 9425 stoves